WARNING: This web is not used anymore. Please use PDBService.MinuteS3May07 instead!
 

Minutes 03 May 07

  • Phone:
    • FNAL: Eric
    • BNL: Carlos, Iris
    • CNAF: Barbara
    • GridKa: Andreas, Doris
    • SARA: Alexander
    • Triumf: Jim Li
    • ATLAS: Sasha, David Front, Richard

  • CERN:
    • 3D: Dirk, Eva, Zbigniew, Gianni
    • POOL: Ioannis
    • ATLAS: Gancho, Florbela
    • LHCb: Marco
    • Taiwan: Dave

  • General announcements:
    • NDGF has already sent the list of database services (see last minutes): Please send a list of the Oracle deployments at the sites to Dirk.
    • Network problems at CERN: protocol negotiation. Eva working on it together with the network people.
    • CHEP sessions on Vancouver. Dirk asks for people who plans to go. There is a BOF session planned a part of the CHEP sessions.
    • Backup and recovery exercise: Barbara offers hosted it at CNAF. Date proposed: 12 and 13 of June. Agenda to be followed up with Barbara. Please let us know of any inconvenience with the date.
    • Availability of databases on wiki. Zbigniew has developed a prototype to introduce the planned interventions directely by the sites.
    • Tickets through GGUS: connection problems tickets disabled. During next weeks we should check if the assignment is correct. If not, resend to ROC so it can be redirected. Tickets latency is about a couple of hours.

  • Sites status
    • Tier 0 network problems - see above
    • Triumf iperf tests in order to find and fix the network problems
    • CNAF: intervention 14 may – maintenance
    • BNL: running smooth, working on the monitoring of the cluster: ganglia, nagios. GGUS ticket latency was 3 hours
    • SARA: one of nodes went down during the morning: full file system. Propagation is reporting a problem with the listener on instance number 1 and cannot connect to the destination queue. Eva just moved the queue and the apply to the available node.
    • GridKa: no interventions, ORA-600 error on the database, Andreas will look on metalink.
    • Taiwan: test with the backup and recovery, working on monitoring

  • Experiment status
    • ATLAS:
      • Preparation to add BNL to the production setup ongoing. We will wait until the network problems are fixed.
      • NDGF was tested during 24 hours very successfully.
      • Scalability test will start end of May – new release of the sw. Client boxes already arranged. * Readback jobs running by Richard: some problems with the queries – being analyzed. Optimizator seems to not pick the right execution plan: full index/table scans. General application problem, not related to the destination sites.
    • LHCB:
      • Waiting for PIC to be joined. Eva has already requested access to PIC LHCb database from the downstream database, but Luis is on holidays.
      • Sw to submit jobs from the grid is ready

  • Stress testing of Tier 1 replicas with ATHENA
    • David Front’s presentation – slides available on the agenda.

  • Frontier/CMS
    • No news from the CMS side.
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-07 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback